40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 1 18:48:13 2011, Jamie, Update, SUS, ETMY sus controller found to be in a bad state 
    Reply  Fri Jul 1 20:22:24 2011, rana, Update, SUS, ETMY sus controller found to be in a bad state 
Message ID: 4933     Entry time: Fri Jul 1 20:22:24 2011     In reply to: 4931
Author: rana 
Type: Update 
Category: SUS 
Subject: ETMY sus controller found to be in a bad state 

Actually, ETMY was the only good one. They should all have the 30 Hz High pass as the damping filter. I think these details are in the elog entry that we originally made while doing ETMY.

They should all also have a 3:30 in the XXSEN to compensate the whitening. The logic is supposed to be that FM1 is ON when the hardware whitening is ON. This is the opposite of the old logic and its why the damping filter has to be moved from 3 to 30 Hz.

ELOG V3.1.3-